# The Super-Kamiokande Collaboration

### v1

@@ F. Nova
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
@@ C. Jesús-Valls
Tokyo Institute of Technology, Department of Physics, Tokyo, Japan
@@ A. Minamino
Hanoi University of Science, Vietnam National University, Hanoi, Vietnam
@@ H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
@@
Tokyo Institute of Technology, Department of Physics, Tokyo, Japan
@@ A.J. Carter
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
@@ V. Kiseeva
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
@@ V. Kurochka
Tokyo Institute of Technology, Department of Physics, Tokyo, Japan
@@ A. Lopez Moreno
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
@@ D.G.R. Martin
Tokyo Institute of Technology, Department of Physics, Tokyo, Japan
@@ D.T. Nguyen
Hanoi University of Science, Vietnam National University, Hanoi, Vietnam
@@ V.V. Tereshchenko
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
@@ H.T. Wallace
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
@@ Abstract
The Super-Kamiokande and T2K collaborations present a joint measurement of neutrino oscillation parameters from their atmospheric and beam neutrino data. It uses a common interaction model for events overlapping in neutrino energy and correlated detector systematic uncertainties between the two datasets, which are found to be compatible. Using 3244.4 days of atmospheric data and a beam exposure of protons on target in (anti)neutrino mode, the analysis finds a 1.9 exclusion of CP-conservation (defined as ) and a preference for the normal mass ordering.
Introduction—Following the observation of neutrino oscillations [1], experiments now aim to fully characterize the three-flavor mixing paradigm described by the Pontecorvo–Maki–Nakagawa–Sakata matrix. Here, neutrino mixing is governed by three mixing angles (, , and ), two mass splittings ( and ), and one Charge Parity (CP) violating phase (). While some oscillation parameters have been precisely measured [2], others remain relatively unconstrained. In particular, the CP-violating phase, the ordering of the neutrino mass states (MO), and the octant of have not been determined experimentally. The Jarlskog invariant [3, 4],
, gives the magnitude of a potential CP violation.
@@
The Tokai-to-Kamioka (T2K) long-baseline neutrino experiment [6] measures neutrino oscillations over a baseline of 295 km using a primarily muon (anti)neutrino beam produced by the neutrino facility at J-PARC, located in Ibaraki, Japan. SK is T2K’s far detector and measures neutrinos after oscillations 2.5∘ off of the beam axis. The beam neutrino flux and neutrino interaction cross sections are constrained by a suite of near detectors (T2K ND) situated 280 m downstream of the neutrino production target.
Motivation for a joint analysis—T2K’s off-axis neutrino beam provides a narrow energy spectrum peaked at 600 MeV and a known direction for beam-induced events at SK. This enables a precise determination of and using the “disappearance” channels (), and some constraint on the CP-violating phase and the MO via the lower statistics “appearance” channels (). The beam composition can be changed from primarily neutrinos to primarily antineutrinos by switching the polarity of electromagnetic focusing horns, providing the ability to compare neutrino and antineutrino oscillations.
T2K’s baseline probes the first oscillation maximum and matter effects have a limited impact. Nonetheless, there are notable parameter degeneracies between the MO and [7]. Degeneracies between the lower and upper octants of also affect the measurements.
@@
SK’s atmospheric neutrino sample, on the other hand, provides a comparatively weak constraint on the atmospheric mixing parameters, due to limited information about the incoming neutrino direction, and a broader range of neutrino energies.
However, upward-going neutrinos experience large matter effects, producing asymmetric oscillations between neutrinos and antineutrinos and providing sensitivity to both the octant of and the MO through the appearance channel at 210 GeV.
This provides a means of breaking the MO– degeneracy, complementing the MO sensitivity achieved at T2K through its lower energy appearance events.
Additionally, common systematic uncertainties at SK can be better constrained in a joint fit than in the individual experiments, due to the addition of the T2K ND constraints to the SK analysis and the large statistics atmospheric sample to the T2K analysis.
@@
Analysis strategy—The analysis described here is based on previous analyses from the two experiments [8, 9], modified to produce a coherent joint analysis. Neutrino oscillation parameters are measured by comparing predictions for the rates and spectra of the atmospheric and beam neutrinos to observations performed at SK and making statistical inferences, following both frequentist and Bayesian approaches. The predictions are made using a model of the two experiments, covering neutrino fluxes, interactions, and detector response, with associated uncertainties. This model is built by unifying aspects of the two experiments’ analyses where appropriate, and using each individual experiment’s approach otherwise.
@@
Due to the similarities in neutrino energy spectra and event selections, T2K and low energy atmospheric events are described by a common neutrino interaction model.
The remaining SK atmospheric events correspond to neutrinos of higher energies, for which the constraint on interaction uncertainties coming from the lower energy neutrino measurements at the T2K ND is not always directly applicable.
As a result, for these events, the same base interaction model is used for the initial simulation, but largely independent parameters are used for interaction systematic uncertainties.
@@
The neutrino flux models of the two experiments [8, 10, 11] are mostly independent, with the only common source of systematic uncertainty coming from hadron production in proton collisions. Hadron production is tuned using different measurements in the two models: the
@@
SK atmospheric flux model uses atmospheric muon measurements [12, 13], whereas T2K’s flux model uses measurements by the NA61/SHINE experiment [14]. The resulting uncertainties on the flux prediction are therefore considered to be independent between the two models.
@@
This topology primarily selects charged current quasi-elastic (CCQE)-like events, although in neutrino running mode an additional sample probes events containing a below-Cherenkov-threshold by requiring exactly one -like ring and one Michel electron.
@@
The T2K beam samples and FC single-ring sub-GeV atmospheric samples in SK have a large kinematic overlap, but differ slightly in their respective event selections, which remain the same as in the publications above.
One additional selection criterion, however, is applied to both the SK FC and T2K samples to remove neutron contamination from the Michel candidates for each event, which changes the event rates by .
Interaction model—Neutrino interactions are simulated with the NEUT generator v5.4.0 [16] using the same configuration as T2K’s analysis [8]. The common “low-energy” uncertainty model used for the T2K and atmospheric SK FC sub-GeV samples is based on T2K’s model, with additions to cover important uncertainties for the SK atmospheric analysis. Additional normalization uncertainties on the neutral current single model are introduced, which scale the resonant and coherent contributions separately, motivated by studies of MiniBooNE data [17, 18].
@@
The ability of the low energy model to describe the atmospheric sub-GeV data samples is evaluated by comparing its predictions using T2K’s ND constraint to the observed data for the down-going part of those samples. The down-going events are mostly unaffected by oscillations, and can therefore be used to test the model without biasing the oscillation measurement.
Good agreement with data is found for the samples targeting events without pions in the final state. For the samples targeting charged current single charged pion (CC1) events, a significant data excess is seen: 225 events are observed for 160.0 12.6(stat) 14.3(syst) predicted in the -like sample, and 84 are observed in the -like sample for 52.0 7.2(stat) 5.3(syst) predicted. The excess is localized at low lepton momentum in the -like sample, and uniformly distributed in the -like sample. An excess is also seen at low momentum in the corresponding -like beam sample, but is not significant due to the low statistics of this sample.
@@
The model for the remaining SK samples (“high-energy model”) is based on that used in the SK analysis, but shares the CCQE part with the low-energy model with the exception of the high (four-momentum transfer) parameters, due to similarities in the CCQE events’ phase space outside of the high region.
SK and T2K’s uncertainty models for resonant production and deep inelastic scattering events are very similar, the main difference being the existence of additional parameters and different nominal values in SK’s analysis. For consistency, the single pion production parameters are set to the same values as in T2K, which come from studies of external data. This increases uncertainties compared to SK’s analysis. The model for pion final-state interactions is tuned to external data [25], as done by T2K.
Data from T2K’s ND are used to constrain uncertainties on the interaction parameters from the low-energy model and the correlated part of the high-energy model, but not on the new uncertainties added for this analysis, or any other aspect of the high-energy model, due to the lack of overlapping phase space between the near-detector and non-sub-GeV atmospheric samples.
Detector model—The same version of the SK event reconstruction algorithm [9] is used for the two experiments, to be able to derive correlations between their detector and reconstruction systematic uncertainties. Many of the detector uncertainties in both SK and T2K’s analyses are estimated from comparisons between atmospheric data and simulation. For these, correlated uncertainties are constructed by evaluating the effects of variations of the detector parameters of the SK model on the number of events in the different samples of the two experiments simultaneously.
Correlations between the reconstructed momentum scale uncertainty of the two experiments are found to have an impact on the constraint on obtained in the data fit. Three of the four analyses described in the next section treat the corresponding parameters as correlated, while the fourth considers them uncorrelated.
The remaining detector uncertainties from the reference SK and T2K analyses that are relevant for only one of the experiments are applied to the corresponding samples.
@@
It allows single ring single Michel electron events with low lepton momentum to migrate between -like and -like samples. The size of this uncertainty covers the excess in data observed for the down-going CC -like events at low momentum.
Oscillation analysis—Two Bayesian and two frequentist analyses based on the model described above but with differences in technical implementation, binning, and statistical methodology are used to fit the data.
To constrain uncertainties using the T2K ND, one analysis performs simultaneous fits of the T2K ND, T2K far detector and SK atmospheric data, while the others use a covariance matrix to encode T2K’s ND constraint on systematic parameters when fitting the events observed at SK.
The Bayesian analyses use Markov Chain Monte Carlo methods to evaluate marginal likelihoods for the parameters of interest, while the frequentist analyses compute the profile likelihood on a fixed grid of values of the oscillation parameters of interest.
@@
For atmospheric oscillation probability calculations, path-dependent density averaging of the matter effect based on a four layer approximation of the PREM model [26] is used, and the fast atmospheric oscillations at low energy are smeared using different techniques depending on the analysis.
The path-dependence yields more precise oscillation probabilities than the conventional approximation assuming layers of constant density.
All analyses utilize a binned log-likelihood test statistic assuming Poisson-distributed bin contents plus Gaussian penalties for the systematic parameters. Finally, reactor experiment measurements of using disappearance, [2, 27, 28, 29, 30], are used as an external constraint.
@@
Robustness studies—Simulated datasets [8] are used to test the robustness of the analysis to effects not included in the base systematic uncertainty model. They are generated using alternative models, and fitted using the nominal model to measure how the obtained -values and constraints on oscillation parameters would be affected by having assumed the wrong model.
@@
Fourteen such simulated datasets are considered, corresponding to alternative neutrino interaction models and data-driven effects at both T2K ND and SK.
@@
These studies are used to estimate, amongst other things, how much the data excess observed in the atmospheric down-going CC samples could bias the results despite the newly added uncertainties if it was coming from an unknown systematic effect. Some of the simulated datasets produce a visible shift in the preferred values for . The uncertainty on is inflated by to account for these effects.
@@
Dataset— The atmospheric dataset is slightly increased compared to [9] to include the full Super-Kamiokande IV period (2008–2018), corresponding to a total live-time of 3244.4 days. The same T2K dataset as in [8] is used, corresponding to exposures of and protons on target in neutrino and antineutrino running modes, respectively.
@@
Bayesian results—The Bayesian analyses assume uniform priors on or , , , and the MO. They find a preference for the normal ordering, and a weak preference for the upper octant (Table 1).
@@
SK and T2K data prefer different octants which leads the joint analysis to have a weaker octant constraint than the individual experiments and a higher probability for maximal mixing, while for the CP-violating phase, both experiments favor similar values (Figure 1).
@@
The exclusion of the CP conserving values of (Figure 2) and is quantified by determining the largest fraction of the posterior density for which the value considered is not included in either of the corresponding highest posterior density credible intervals obtained by the two Bayesian analyses (Table 2).
@@
Figure 2: The posterior density for the Jarlskog invariant with credible intervals overlaid, marginalized over both MOs, and assuming a uniform prior in either or .
Frequentist results—Ensembles of pseudo-experiments are constructed
to evaluate the frequentist significance of the CP and MO results, taking into account statistical fluctuations and randomizing the values of nuisance oscillation and systematic parameters according to their posterior [31] and prior probability distributions respectively.
Estimating the significance of CP conservation (CPC) based on the presence or not of both and in the confidence intervals was found to have significant over-coverage.
@@
The inverted ordering is disfavored while good agreement with the normal ordering hypothesis is found, resulting in a parameter [32] for the inverted ordering of 0.18. The distribution of the MO test statistic depends on the assumed values of (from SK samples) and (from T2K samples). The -value for the inverted ordering varies between 0.05 and 0.08 when assuming different fixed true values for and over the range of their 90% confidence intervals. The best-fit values and 68.3% confidence intervals obtained using the Feldman–Cousins method [33] where necessary, are ,
@@
Frequentist -values for the different CP and MO hypotheses.
@@
The frequentist -values [35] additionally show good consistency between the values of the systematic parameters favored by the T2K ND and atmospheric data (), as well as between the atmospheric and beam samples ().
Discussion—The SK and T2K datasets favor similar values for the CP phase, close to maximal CP violation, and both show a preference for the normal MO. As a result, the combined analysis finds increased preferences for non-conservation of CP symmetry and the normal ordering.
When looking directly at the exclusion of CPC through the presence of in credible intervals or frequentist -values, an exclusion at the 2 level is found. However, the significance can fall below 2 due to the potential weaknesses of the uncertainty model tested using simulated datasets. The alternative models having the most impact are the one assuming that the excess seen for down-going CC1-like events is fully due to an unknown systematic effect, and an alternative nuclear model for CCQE interactions [36]. This indicates areas where the robustness of the model should be improved for future analyses.
Conclusion—The SK and T2K collaborations have produced a first joint analysis of their data. A common neutrino interaction and detector model has been developed for the events of the two experiments which overlap in energy, and is found to be able to properly describe the two datasets. The results show an exclusion of the CP-conserving value of the Jarlskog invariant with a significance between 1.9 and 2.0, a limited preference for the normal ordering, and no strong preference for the octant. This first joint analysis is also an important step towards the combined beam and atmospheric data analyses planned by next generation neutrino oscillation experiments.

### v2

@@ F. Nova
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
@@ C. Jesús-Valls
Institute of Science Tokyo, Department of Physics, Tokyo, Japan
@@ A. Minamino
VNU University of Science, Vietnam National University, Hanoi, Vietnam
@@ H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
@@
Institute of Science Tokyo, Department of Physics, Tokyo, Japan
@@ A.J. Carter
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
@@ V. Kiseeva
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
@@ V. Kurochka
Institute of Science Tokyo, Department of Physics, Tokyo, Japan
@@ A. Lopez Moreno
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
@@ D.G.R. Martin
Institute of Science Tokyo, Department of Physics, Tokyo, Japan
@@ D.T. Nguyen
VNU University of Science, Vietnam National University, Hanoi, Vietnam
@@ V.V. Tereshchenko
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
@@ H.T. Wallace
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
@@ Abstract
The Super-Kamiokande and T2K collaborations present a joint measurement of neutrino oscillation parameters from their atmospheric and beam neutrino data. It uses a common interaction model for events overlapping in neutrino energy and correlated detector systematic uncertainties between the two datasets, which are found to be compatible. Using 3244.4 days of atmospheric data and a beam exposure of protons on target in (anti)neutrino mode, the analysis finds a 1.9 exclusion of CP-conservation (defined as ) and a 1.2 exclusion of the inverted mass ordering.
Introduction—Following the observation of neutrino oscillations [1], experiments now aim to fully characterize the three-flavor mixing paradigm described by the Pontecorvo–Maki–Nakagawa–Sakata matrix using neutrinos from different sources [2, 3, 4, 5, 6, 7]. Here, neutrino mixing is governed by three mixing angles (, , and ), two mass splittings ( and ), and one Charge Parity (CP) violating phase (). While some oscillation parameters have been precisely measured [8], others remain relatively unconstrained. In particular, the CP-violating phase, the ordering of the neutrino mass states (MO), and the octant of have not been determined experimentally.
The magnitude of CP violation is proportional to the Jarlskog invariant [9, 10, 11],
@@
The Tokai-to-Kamioka (T2K) long-baseline neutrino experiment [13] measures neutrino oscillations over a baseline of 295 km using a primarily muon (anti)neutrino beam produced by the neutrino facility at J-PARC, located in Ibaraki, Japan. SK is T2K’s far detector (FD) and measures neutrinos after oscillations 2.5∘ off of the beam axis. The beam neutrino flux and neutrino interaction cross sections are constrained by a suite of near detectors (T2K ND) situated 280 m downstream of the neutrino production target.
Motivation for a joint analysis—T2K and SK have complementary strengths to study neutrino oscillations. T2K’s off-axis neutrino beam provides a narrow energy spectrum peaked at 600 MeV and a known direction for beam-induced events at SK. This enables a precise measurement of the “disappearance” of through oscillations, which manifest as a dip around 600 MeV in the spectra of events observed at SK.
T2K precisely measures and , which are connected to the peak energy and the amplitude of this disappearance, respectively. The MO and the value of both affect the “appearance” probabilities for neutrinos P() and antineutrinos P() asymmetrically.
T2K’s ability to change the beam composition from primarily neutrinos to primarily antineutrinos gives a powerful way to compare the oscillations of the two, enabling some constraint on and the MO from the numbers of and events observed at the far detector.
The similarity of the effects of the MO and the value of can however seriously degrade the experiment’s ability to measure these parameters (Appendix A).
@@
SK’s atmospheric neutrino sample, on the other hand, provides a comparatively weak constraint on and , due to limited information about the incoming neutrino direction, and a broader range of neutrino energies. However, upward-going neutrinos experience large matter effects, which asymmetrically modify the oscillation probabilities of neutrinos and antineutrinos, dependent upon the MO.
In particular, 210 GeV (anti-)neutrinos experience a resonant enhancement of their appearance probability if the ordering is normal (inverted). The size of this enhancement is proportional to . This provides sensitivity to both the octant of and the MO through an excess of upward-going or events at these energies. In a combined analysis, this provides a means of breaking the T2K MO– degeneracy, complementing the MO sensitivity achieved at T2K through its lower energy appearance events.
Additionally, systematic uncertainties at SK are better constrained in a joint fit than in the individual experiments: the beam and atmospheric samples at SK share numerous detector uncertainties and both can receive neutrino interaction uncertainty constraints from the T2K ND.
@@
Analysis strategy—The analysis described here is based on previous analyses from the two experiments [3, 14], modified to produce a coherent joint analysis. Neutrino oscillation parameters are measured by comparing predictions for the rates and spectra of the atmospheric and beam neutrinos to observations performed at SK. The predictions are made using a model of the two experiments, covering neutrino fluxes, interactions, and detector response, with associated uncertainties. This model is built unifying aspects of the two experiments’ analyses where appropriate, and using each individual experiment’s approach otherwise.
@@
Due to similarities in the neutrino energy spectra and event selections, T2K and low energy atmospheric events are described by a common neutrino interaction model. Largely independent interaction systematic uncertainties are used for higher energy atmospheric events as measurements from the T2K ND are not always applicable to these events, though the base interaction model is the same for all events.
@@
The neutrino flux models for the experiments [3, 15, 16] are mostly independent, with the only common source of systematic uncertainty coming from hadron production in proton collisions.
Hadron production is tuned using different, independent measurements in the two models: the
@@
SK atmospheric flux model uses atmospheric muon measurements [17, 18], whereas T2K’s model uses measurements by the NA61/SHINE experiment [19].
@@
This topology primarily selects charged current quasi-elastic (CCQE)-like events.
However, in neutrino running mode an additional sample probes events containing a below-Cherenkov-threshold by requiring exactly one -like ring and one Michel electron.
@@
The T2K beam samples and FC single-ring sub-GeV atmospheric samples in SK have a large kinematic overlap, but differ slightly in their respective event selections.
The selections remain unchanged relative to the publications above.
One additional selection criterion, however, is applied to all SK FC and T2K samples to remove neutron contamination from the Michel candidates for each event.
This cut changes the event rates by .
Interaction model—Neutrino interactions are simulated with the NEUT generator v5.4.0 [21] using the same configuration as T2K’s analysis [3]. The common “low-energy” uncertainty model used for the T2K and atmospheric SK FC sub-GeV samples is based on T2K’s model, with two additions to cover important uncertainties for the atmospheric samples. Additional normalization uncertainties on the neutral current single model are introduced
motivated by studies of MiniBooNE data [22, 23].
These uncertainties separately scale the resonant and coherent components.
@@
This uncertainty changes the number of events in the e-like atmospheric sub-GeV sample targeting events without pions in the final state (CC0) by 2.2%, and is the interaction uncertainty with the largest impact on the atmospheric sample contribution to the measurement.
The ability of the low energy model to describe the atmospheric sub-GeV samples is evaluated by comparing its predictions to the observed down-going data.
Those events are mostly unaffected by oscillations and can therefore be used to test the model without biasing the oscillation measurement.
Good agreement with data is found for the samples targeting CC0 events when using the T2K ND constraint, while a small data excess is seen without it.
For the samples targeting charged current single charged pion (CC1) events, a significant data excess is seen: 225 events are observed in the -like sample for 160.0 12.6(stat) 14.3(syst) predicted using the T2K ND constraint and 84 events are observed in the -like sample for 52.0 7.2(stat) 5.3(syst) predicted.
The T2K ND constraint reduces prediction in these samples by 20% compared to the nominal prediction.
The excess is localized at low lepton momentum in the -like sample, and uniformly distributed in the -like sample.
An excess is also seen at low momentum in the corresponding -like beam sample, but is not significant due to the low statistics of this sample.
@@
The model for the remaining SK samples (“high-energy model”) is based on that used in the SK analysis, with its pion secondary interaction and CCQE parts shared with the low-energy model.
However, two high (four-momentum transfer) CCQE normalization parameters are left uncorrelated with the low-energy model due to limited phase space overlap.
The high energy model for pion final-state interactions is tuned to external data [30], as is done by T2K.
Due to little overlap between the phase spaces of the near-detector and non-sub-GeV atmospheric samples, data from T2K’s ND are used to constrain interaction uncertainties in the low-energy model and the correlated part of the high-energy model, excluding newly added uncertainties and other parts of the high-energy model.
Detector model—
Many of the detector uncertainties in the SK and T2K analyses are estimated from comparisons between atmospheric data and simulation.
For these, correlated uncertainties are constructed by simultaneously evaluating the effects of detector parameter variations on the event rate in both the SK and T2K samples.
Correlations between the reconstructed momentum scale uncertainties of the two experiments are found to have an impact on the constraint obtained in the data fit.
Other detector uncertainties from the reference analyses that are relevant for only one of the experiments are applied to the corresponding samples here.
@@
It allows single-ring single-Michel electron events with low lepton momentum to migrate between the -like and -like samples. The size of this uncertainty covers the excess in data observed for the down-going CC -like events at low momentum.
Oscillation analysis—Two Bayesian and two frequentist analyses were constructed (Appendix C).
When two of those frameworks return different results for a given measurement, the more conservative option is reported.
@@
For atmospheric oscillation probability calculations, path-dependent density averaging of matter effects based on a four layer approximation of the PREM model [31] is used and fast atmospheric oscillations at low energy are smeared. The path-dependence yields more precise oscillation probabilities than the conventional approximation assuming layers of constant density.
Reactor experiment measurements of using disappearance, [8, 4, 5, 32, 33], are used as an external constraint.
@@
Robustness studies—Simulated datasets [3], generated using alternative models and fit using the nominal model are used to measure how -values and oscillation parameter constraints would be affected if the assumed model is incomplete (Appendix D).
@@
Fourteen simulated datasets are considered, corresponding to alternative neutrino interaction models and data-driven effects at both T2K ND and SK.
@@
These studies are used to estimate, for example, how the observed atmospheric down-going CC data excess could bias the results if it originated from an unknown systematic effect.
Some of the simulated datasets produce a visible shift in the preferred values for .
The uncertainty on is therefore inflated by to account for these effects.
@@
Dataset— The atmospheric dataset is slightly increased compared to Ref. [14] to include the full Super-Kamiokande IV period (2008–2018), corresponding to a total live-time of 3244.4 days. The same T2K dataset as Ref. [3] is used, corresponding to exposures of and protons on target in neutrino and antineutrino modes, respectively.
@@
Bayesian results—The Bayesian analyses assume uniform priors on or , , , and the MO. They find a preference for the normal ordering and a weak preference for the upper octant (Table 1).
@@
SK and T2K data prefer different octants, which leads the joint analysis to have a weaker octant constraint than the individual experiments and a stronger preference for maximal mixing.
Both experiments favor similar values of the CP-violating phase (Figure 1).
@@
The exclusion of CP-conserving values of (Figure 2) and is reported as the largest fraction of the posterior density for which that value is not included in either of the two Bayesian analyses’ highest posterior density credible intervals
(Table 2).
@@
Figure 2: Posterior density for the Jarlskog invariant with credible intervals overlaid, marginalized over both MOs, and assuming a uniform prior in either or .
The left edges of the intervals are close to each other in the region .
Frequentist results— The frequentist significance of the CP and MO results is evaluated using ensembles of pseudo-experiments. Estimating the significance of CP conservation (CPC) based on the presence or absence of both and in the confidence intervals was found to have significant over-coverage.
@@
The inverted ordering is disfavored while good agreement with the normal ordering hypothesis is found, resulting in a parameter [34] for the inverted ordering of 0.18.
The best-fit values and 68.3% confidence intervals obtained using the Feldman–Cousins method [35] where necessary,
are ,
@@
Frequentist -values for different CP and MO hypotheses.
@@
The frequentist -values [37] additionally show consistency between the values of the systematic parameters favored by the T2K ND and atmospheric data (), as well as between the atmospheric and beam samples ().
Discussion—The SK and T2K datasets favor similar values for the CP phase, close to maximal CP violation, and both show a preference for the normal MO. As a result, the combined analysis finds increased preferences for CP non-conservation and the normal ordering.
When looking directly at the exclusion of CPC through the presence of in credible intervals or frequentist -values, an exclusion at the 2 level is found.
However, the significance can fall below 2 due to potential weaknesses of the uncertainty model tested using simulated datasets.
The alternative model assuming that the down-going CC1-like data excess is completely due to an unknown systematic effect and the alternative nuclear model for CCQE interactions [38]
have the largest impact.
The -value obtained for the inverted ordering is significantly larger than the one obtained in the reference atmospheric analysis () [14], despite the addition of the T2K samples which also favor the normal ordering.
This was found to come mostly from the different values of favored by the two analyses (Appendix E).
The constraint from the T2K ND has a small effect on the measurement of and the MO using atmospheric samples. The most important systematic uncertainties for this measurement are related to the atmospheric neutrino flux below 1 GeV, the cross-section ratio and the particle identification for single ring events, which cannot be directly constrained using the T2K ND. For the MO, most of the atmospheric neutrino sensitivity comes from the high energy samples where this constraint is not used for most interaction modes.
Accordingly, the MO sensitivity benefits mainly from the stronger
constraint provided by T2K data.
Conclusion—The SK and T2K collaborations have produced a first joint analysis of their data. Common neutrino interaction and detector models have been developed for events from the two experiments with overlapping energies and are found to properly describe both datasets.
The results show an exclusion of the CP-conserving value of the Jarlskog invariant with a significance between 1.9 and 2.0, a limited preference for the normal ordering with a 1.2 exclusion of the inverted ordering [39], and no strong preference for the octant. This first joint analysis is an important step towards the combined beam and atmospheric data analyses planned by next-generation neutrino oscillation experiments.
The data related to this work can be found in Ref. [40].